OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3. y - 2 = -3(x + 7) y + 7 = -3(x - 2) y + 2 = -3(x - 7) y - 7 = -3(x + 2)

OpenStudy (angela210793):

y + 2 = -3(x - 7)

OpenStudy (katrinakaif):

(y + 2) = -3(x - 7)

OpenStudy (angela210793):

y-y1=m(x-x1)